OpenStudy (anonymous):

What is the solution of the equation? 36x^2 = 100 -9/25 , 9/25 -3/5 , 3/5 -25/9 , 25/9 -5/3 , 5/3

OpenStudy (anonymous):

36x^2=100 x^2=25/9 Can you solve the rest?

OpenStudy (anonymous):

Remember a square root has a positive and negative answer.

OpenStudy (anonymous):

x^2 = 100/36 (divide by 36 on both sides) x^2 = 25/9 (reduce fraction) x = sqrt(25/9) (square root both sides) x = 5/3 (simplify. the sqrt of 25 = 5 and the sqrt of 9 = 3)

OpenStudy (anonymous):

Don't just type the answer next time :P .

OpenStudy (anonymous):

@QuantumTorch : A square root is positive and negative. So it's -5/3 AND 5/3

OpenStudy (anonymous):

thank yall sooo much .

OpenStudy (anonymous):

Welcome :) .

OpenStudy (anonymous):

I didn't put both +/- because they were already in the answers, so I figured that was assumed. :P

OpenStudy (anonymous):

36x^2 = 100 x^2 = 100/36 x = +/- (10/6) 1. What is the solution of the equation? 36x2 = 100 (1 point) (0 pts) –9/25, 9/25 (0 pts) –3/5, 3/5 (0 pts) –25/9, 25/9 (1 pt) –5/3, 5/3 1 /1 point

OpenStudy (anonymous):

whats the rest ?? @whiterabbit541